OpenStudy (anonymous):

Trying to understand how sqrt(27x^8) becomes 3^3/2 X^4

OpenStudy (anonymous):

the square root is just like raising to the power of 1/2. so sqrt(27x^8) becomes (27x^8)^1/2 27 = 3^3, so: (3^3 * x^8)^1/2 and then for a power raised to a power, you multiply the exponents: 3^(3*1/2) * x^ (8*1/2) and get 3^(3/2) * x^4

jimthompson5910 (jim_thompson5910):

sqrt(27) = sqrt(9*3) = sqrt(9)*sqrt(3) = 3*sqrt(3) sqrt(x^8) =sqrt((x^4)^2) = x^4 So the answer is 3x^4*sqrt(3)

OpenStudy (anonymous):

and \[3 \sqrt{3} = 3 \times 3^{\frac{1}{2}} = 3^{1+\frac{1}{2}} = 3^{\frac{3}{2}}\]